# A relative version of Bass’ theorem about finite-dimensional algebras

### v1

@@ Introduction
According to the orginal Bass’ definition in [1], an associative ring $A$ is said to be left perfect if all left $A$-modules have projective covers. By [1, Theorem P(3)], the ring $A$ is left perfect if and only if all flat left $A$-modules are projective. The condition of [1, Theorem P(1)] provides a characterization of left perfect rings in terms of their structural properties that is used in the argument above.
@@
The thematic example of a cotorsion pair if the flat cotorsion pair (also known as the Enochs cotorsion pair [8], [9, Definition 5.18(i)]). In this example, the class $\mathsf{F}=A{\operatorname{\mathsf{--Mod}}}_{\mathsf{flat}}$ is the class of all flat left $A$-modules. The class $\mathsf{C}=A{\operatorname{\mathsf{--Mod}}}^{\mathsf{cot}}$ is the class of all cotorsion left $A$-modules. Here a left $A$-module $C$ is said to be cotorsion if $\operatorname{Ext}^{1}_{A}(F,C)=0$ for all flat left $A$-modules $F$, or equivalently, $\operatorname{Ext}^{n}_{A}(F,C)=0$ for all flat left $A$-modules $F$ and all $n\geq 1$.
The main result about cotorsion pairs is the Eklof–Trlifaj theorem [6, Theorems 2 and 10], [9, Theorem 6.11 and Corollary 6.14]. The Eklof–Trlifaj theorem concerns the notion of a complete cotorsion pair, and its simplest formulation claims that every cotorsion pair generated by a set of modules is complete. Proving completeness of the flat cotorsion pair in the category of modules over an arbitrary ring was a major open problem, which became known under the name of the Flat Cover Conjecture. Posed by Enochs in [7] in 1981 (see also [16, Section 1.3]), it was proved by Bican–El Bashir–Enochs in [3] in 2001. One of the two proofs of the Flat Cover Conjecture presented in [3] was based on the Eklof–Trlifaj theorem.
@@
Specifically, the flaprojective cotorsion pair $(\mathsf{F},\mathsf{C})$ is formed by the following two classes of left $A$-modules. The right-hand class $\mathsf{C}$ consists of all left $A$-modules that are cotorsion as $R$-modules, i. e., the $A$-modules whose underlying $R$-module is cotorsion. The left-hand class $\mathsf{F}$ consists of all the left $A$-modules $F$ such that $\operatorname{Ext}^{1}_{A}(F,C)=0$ for all left $A$-modules $C$ that are cotorsion over $R$. The $A$-modules $F$ satisfying the latter condition are called flaprojective relative to $R$ (or $A/R$-flaprojective) [13, Section 1.6].
The definition of a flaprojective module in the previous paragraph is very indirect and roundabout. First one defines cotorsion $R$-modules as the $R$-modules right $\operatorname{Ext}^{1}_{R}$-orthogonal of flat $R$-modules, and then one defined flaprojective $A$-modules as the $A$-modules left $\operatorname{Ext}^{1}_{A}$-orthogonal to $R$-cotorsion $A$-modules. A much more direct and explicit description is provided by the Eklof–Trlifaj theorem, but it uses the concept of an ordinal-indexed filtration on modules. A left $A$-module $F$ is $A/R$-flaprojective if and only if $F$ is a direct summand of an $A$-module admitting an ordinal-indexed filtration whose successive quotients are the $A$-modules $A\otimes_{R}G$ induced from flat left $R$-modules $G$.
@@
There are many examples of left perfect rings and left perfect $k$-algebras beyond the finite-dimensional $k$-algebras. Are there any nontrivial classes of examples of relatively left perfect ring homomorphisms beyond the ones described by Theorem 0.1 ? Can one characterize such ring homomorphisms by any relative analogues of the properties listed in [1, Theorem P] ?
@@ Acknowledgement
I am grateful to Michal Hrbek and Sergio Pavon for helpful conversations. The author is supported by the GAČR project 23-05148S and the Institute of Mathematics, Czech Academy of Sciences (research plan RVO: 67985840).
@@ 1. Preliminaries on Cotorsion Pairs
A pair of classes $(\mathsf{F},\mathsf{C})$ in $A{\operatorname{\mathsf{--Mod}}}$ is said to be a cotorsion pair if $\mathsf{C}=\mathsf{F}^{\perp_{1}}$ and $\mathsf{F}={}^{\perp_{1}}\mathsf{C}$. For any class of modules $\mathsf{S}\subset A{\operatorname{\mathsf{--Mod}}}$, the pair of classes $\mathsf{F}=\mathsf{S}^{\perp_{1}}$ and $\mathsf{C}={}^{\perp_{1}}\mathsf{C}$ is a cotorsion pair in $A{\operatorname{\mathsf{--Mod}}}$. The latter cotorsion pair $(\mathsf{F},\mathsf{C})$ is said to be generated by the class of modules $\mathsf{S}\subset A{\operatorname{\mathsf{--Mod}}}$.
@@
[…]
with $A$-modules $F$, $F^{\prime}\in\mathsf{F}$ and $C$, $C^{\prime}\in\mathsf{C}$. The short exact sequence (1) is called a special precover sequence. The short exact seqeuence (2) is called a special preenvelope sequence. Taken together, the short exact sequences (12) are called the approximation sequences.
An introductory discussion of cotorsion pairs can be found in the introduction to [12]. Many examples of approximation sequences (for abelian groups, i. e., modules over the ring of integers $A=\mathbb{Z}$) are discussed in [11, Section 12]. The main reference source on cotorsion pairs in module categories is the book [9].
@@
This is [6, Lemma 1] or [9, Lemma 6.2]. ∎
@@
(b) Assume that the free left $A$-module $A$ with one generator belongs to $\mathsf{S}$, that is ${}_{A}A\in\mathsf{S}$. Then one has $\mathsf{F}=\mathsf{Fil}(\mathsf{S})^{\oplus}$.
@@
This is [6, Theorems 2 and 10] or [9, Theorem 6.11 and Corollary 6.14]. ∎
@@ 2. Cotorsion Modules
Using the facts that all projective left $R$-modules are flat and the class of flat modules is closed under kernels of epimorphisms in $R{\operatorname{\mathsf{--Mod}}}$, one can easily show that $\operatorname{Ext}_{R}^{n}(F,C)=0$ for all flat left $R$-modules $F$, all cotorsion left $R$-modules $C$, and all integers $n\geq 1$ (cf. [9, Lemma 5.24] and [14, Lemma 7.1]). It follows that the class of cotorsion $R$-modules is closed under cokernels of monomorphisms in $R{\operatorname{\mathsf{--Mod}}}$; it is also obviously closed under extensions and infinite products. We denote the class of flat left $R$-modules by $R{\operatorname{\mathsf{--Mod}}}_{\mathsf{flat}}\subset R{\operatorname{\mathsf{--Mod}}}$ and the class of cotorsion left $R$-modules by $R{\operatorname{\mathsf{--Mod}}}^{\mathsf{cot}}\subset R{\operatorname{\mathsf{--Mod}}}$.
@@
This is [2, Theorem 5.1(2)]. See also [2, Theorem 1.2(2) or Proposition 4.8(2)] for essentially equivalent formulations of this theorem. A further discussion in category-theoretic context can be found in [14, Corollary 9.2 and Proposition 9.4]. ∎
@@
[…]
@@
This is [10, Corollary 1.6.4]. The point is that any $A$-module can be embedded into a cotorsion $A$-module (in fact, all injective $A$-modules are cotorsion by definition). Therefore, any $A$-module $M$ has a coresolution (right resolution) by cotorsion $A$-modules, i. e., there is an exact complex of $A$-modules
[…]
@@
[…]
@@
This is [10, Lemma 1.3.4(a)] or [13, Lemma 1.6(a)]. One can use the $\operatorname{Ext}^{1}$-adjunction of change of rings as in [5, Proposition VI.4.1.3] (see Lemma 3.4 below). Alternatively, the following elementary argument works. A left $R$-module $C$ is cotorsion if and only if the functor $\operatorname{Hom}_{R}({-},C)$ takes every short exact sequence of flat left $R$-modules to a short exact sequence of abelian groups. Given a short exact sequence of flat left $R$-modules $0\longrightarrow H\longrightarrow G\longrightarrow F\longrightarrow 0$, we have a short exact sequence of flat left $A$-modules $0\longrightarrow A\otimes_{R}H\longrightarrow A\otimes_{R}G\longrightarrow A\otimes_{R}F\longrightarrow 0$. By assumption, the $A$-module $D$ is cotorsion, so the functor $\operatorname{Hom}_{A}({-},D)$ takes the latter short exact sequence of $A$-modules to a short exact sequence of abelian groups. It remain to recall the adjunction isomorphism of abelian groups $\operatorname{Hom}_{A}(A\otimes_{R}M,\>D)\simeq\operatorname{Hom}_{R}(M,D)$, which holds for any left $R$-module $M$. The assertion of the lemma is also a special case of the next Lemma 2.4. ∎
@@
This is [10, Lemma 1.3.3(a)] or [13, Lemma 1.5(a)]. The point is that the tensor product $K\otimes_{S}G$ is a flat left $R$-module for any flat left $S$-module $G$. Based on this observation, one can use the $\operatorname{Ext}^{1}$-adjunction of the tensor product and Hom functors as in [5, formula (4) in Section XVI.4], [10, formula (1.1) in Section 1.2], or [13, Lemma 1.1]. Alternatively, an elementary argument can be found in [13, proof of Lemma 1.5]. ∎
@@
Let $R\longrightarrow A$ be a homomorphism of associative rings making $A$ a finitely generated projective right $R$-module. Consider the $R$-$A$-bimodule
@@
By Lemma 2.5, we have a left $A$-module isomorphism $A\otimes_{R}C\simeq\operatorname{Hom}_{R}(K,C)$, where $K=\operatorname{Hom}_{R^{\mathrm{op}}}(A,R)$. The $R$-$A$-bimodule $K$ is finitely generated and projective, hence flat, as a left $R$-module. So Lemma 2.4 is applicable for the ring $S=A$. ∎
@@
The lemma claims that the complex (3) is acyclic. In fact, the complex (3) is contractible as a complex of left $R$-modules (but not as a complex of left $A$-modules). The natural $R$-linear contracting homotopy $h$ is given by the formula
@@
Proceeding in this way, we prove by induction that the left $A$-module $A^{\otimes_{R}\,n}\otimes_{R}D$ (with the $A$-module structure coming from the left $A$-module structure on the leftmost tensor factor $A$) is cotorsion for all $n\geq 1$. Now Lemma 2.7 provides a left resolution of $D$ by cotorsion $A$-modules. By Corollary 2.2, we can conclude that $D$ is a cotorsion $A$-module, as desired. ∎

### v2

@@ Introduction
According to the original Bass’ definition in [1], an associative ring $A$ is said to be left perfect if all left $A$-modules have projective covers. By [1, Theorem P(3)], the ring $A$ is left perfect if and only if all flat left $A$-modules are projective. The condition of [1, Theorem P(1)] provides a characterization of left perfect rings in terms of their structural properties that is used in the argument above.
@@
The thematic example of a cotorsion pair is the flat cotorsion pair (also known as the Enochs cotorsion pair [8], [10, Definition 5.18(i)]). In this example, the class $\mathsf{F}=A{\operatorname{\mathsf{--Mod}}}_{\mathsf{flat}}$ is the class of all flat left $A$-modules. The class $\mathsf{C}=A{\operatorname{\mathsf{--Mod}}}^{\mathsf{cot}}$ is the class of all cotorsion left $A$-modules. Here a left $A$-module $C$ is said to be cotorsion if $\operatorname{Ext}^{1}_{A}(F,C)=0$ for all flat left $A$-modules $F$, or equivalently, $\operatorname{Ext}^{n}_{A}(F,C)=0$ for all flat left $A$-modules $F$ and all $n\geq 1$.
The main result about cotorsion pairs is the Eklof–Trlifaj theorem [6, Theorems 2 and 10], [10, Theorem 6.11 and Corollary 6.14]. The Eklof–Trlifaj theorem concerns the notion of a complete cotorsion pair, and its simplest formulation claims that every cotorsion pair generated by a set of modules is complete. Proving completeness of the flat cotorsion pair in the category of modules over an arbitrary ring was a major open problem, which became known under the name of the Flat Cover Conjecture. Posed by Enochs in [7] in 1981 (see also [18, Section 1.3]), it was proved by Bican–El Bashir–Enochs in [3] in 2001. One of the two proofs of the Flat Cover Conjecture presented in [3] was based on the Eklof–Trlifaj theorem.
@@
Specifically, the flaprojective cotorsion pair $(\mathsf{F},\mathsf{C})$ is formed by the following two classes of left $A$-modules. The right-hand class $\mathsf{C}$ consists of all left $A$-modules that are cotorsion as $R$-modules, i. e., the $A$-modules whose underlying $R$-module is cotorsion. The left-hand class $\mathsf{F}$ consists of all the left $A$-modules $F$ such that $\operatorname{Ext}^{1}_{A}(F,C)=0$ for all left $A$-modules $C$ that are cotorsion over $R$. The $A$-modules $F$ satisfying the latter condition are called flaprojective relative to $R$ (or $A/R$-flaprojective) [14, Section 1.6].
The definition of a flaprojective module in the previous paragraph is very indirect and roundabout. First one defines cotorsion $R$-modules as the $R$-modules right $\operatorname{Ext}^{1}_{R}$-orthogonal to flat $R$-modules, and then one defines flaprojective $A$-modules as the $A$-modules left $\operatorname{Ext}^{1}_{A}$-orthogonal to $R$-cotorsion AA-modules. A much more direct and explicit description is provided by the Eklof–Trlifaj theorem, but it uses the concept of an ordinal-indexed filtration on modules. A left $A$-module $F$ is $A/R$-flaprojective if and only if $F$ is a direct summand of an $A$-module admitting an ordinal-indexed filtration whose successive quotients are the $A$-modules $A\otimes_{R}G$ induced from flat left $R$-modules $G$.
@@
There are many examples of left perfect rings and left perfect $k$-algebras beyond the finite-dimensional $k$-algebras. A class of examples of relatively left perfect ring homomorphisms is provided Theorem 0.1. Another such class, consisting of all the left flat ring epimorphisms $R\longrightarrow A$, is discussed in Example 3.10 at the end of this paper. What are other nontrivial classes of examples of relatively left perfect ring homomorphisms? Can one characterize relatively left perfect ring homomorphisms by any relative analogues of the properties listed in [1, Theorem P] ?
@@ Acknowledgement
I am grateful to Michal Hrbek and Sergio Pavon for helpful conversations. A special gratitute goes to Silvana Bazzoni for communicating Example 3.10 to me and giving a kind permission to reproduce it in this paper. The author is supported by the GAČR project 23-05148S and the Institute of Mathematics, Czech Academy of Sciences (research plan RVO: 67985840).
@@ 1. Preliminaries on Cotorsion Pairs
A pair of classes $(\mathsf{F},\mathsf{C})$ in $A{\operatorname{\mathsf{--Mod}}}$ is said to be a cotorsion pair if $\mathsf{C}=\mathsf{F}^{\perp_{1}}$ and $\mathsf{F}={}^{\perp_{1}}\mathsf{C}$. For any class of modules $\mathsf{S}\subset A{\operatorname{\mathsf{--Mod}}}$, the pair of classes $\mathsf{C}=\mathsf{S}^{\perp_{1}}$ and $\mathsf{F}={}^{\perp_{1}}\mathsf{C}$ is a cotorsion pair in $A{\operatorname{\mathsf{--Mod}}}$. The latter cotorsion pair $(\mathsf{F},\mathsf{C})$ is said to be generated by the class of modules $\mathsf{S}\subset A{\operatorname{\mathsf{--Mod}}}$.
@@
(1) $0$ $\mskip 1.5mu\relbar\joinrel\relbar\joinrel\rightarrow\mskip 1.5muC^{\prime}$ $\mskip 1.5mu\relbar\joinrel\relbar\joinrel\rightarrow\mskip 1.5muF$ $\mskip 1.5mu\relbar\joinrel\relbar\joinrel\rightarrow\mskip 1.5muM$ $\mskip 1.5mu\relbar\joinrel\relbar\joinrel\rightarrow\mskip 1.5mu0,$
(2) $0$ $\mskip 1.5mu\relbar\joinrel\relbar\joinrel\rightarrow\mskip 1.5muM$ $\mskip 1.5mu\relbar\joinrel\relbar\joinrel\rightarrow\mskip 1.5muC$ $\mskip 1.5mu\relbar\joinrel\relbar\joinrel\rightarrow\mskip 1.5muF^{\prime}$ $\mskip 1.5mu\relbar\joinrel\relbar\joinrel\rightarrow\mskip 1.5mu0$
with $A$-modules $F$, $F^{\prime}\in\mathsf{F}$ and $C$, $C^{\prime}\in\mathsf{C}$. The short exact sequence (1) is called a special precover sequence. The short exact sequence (2) is called a special preenvelope sequence. Taken together, the short exact sequences (12) are called the approximation sequences.
An introductory discussion of cotorsion pairs can be found in the introduction to [13]. Many examples of approximation sequences (for abelian groups, i. e., modules over the ring of integers $A=\mathbb{Z}$) are discussed in [12, Section 12]. The main reference source on cotorsion pairs in module categories is the book [10].
@@
This is [6, Lemma 1] or [10, Lemma 6.2]. ∎
@@
(b) Assume that the free left $A$-module AA with one generator belongs to $\mathsf{S}$, that is $_{A}A\in\mathsf{S}$. Then one has $\mathsf{F}=\mathsf{Fil}(\mathsf{S})^{\oplus}$.
@@
This is [6, Theorems 2 and 10] or [10, Theorem 6.11 and Corollary 6.14]. ∎
@@ 2. Cotorsion Modules
Using the facts that all projective left $R$-modules are flat and the class of flat modules is closed under kernels of epimorphisms in $R{\operatorname{\mathsf{--Mod}}}$, one can easily show that $\operatorname{Ext}_{R}^{n}(F,C)=0$ for all flat left $R$-modules $F$, all cotorsion left $R$-modules $C$, and all integers $n\geq 1$ (cf. [10, Lemma 5.24] and [15, Lemma 7.1]). It follows that the class of cotorsion $R$-modules is closed under cokernels of monomorphisms in $R{\operatorname{\mathsf{--Mod}}}$; it is also obviously closed under extensions and infinite products. We denote the class of flat left $R$-modules by $R{\operatorname{\mathsf{--Mod}}}_{\mathsf{flat}}\subset R{\operatorname{\mathsf{--Mod}}}$ and the class of cotorsion left $R$-modules by $R{\operatorname{\mathsf{--Mod}}}^{\mathsf{cot}}\subset R{\operatorname{\mathsf{--Mod}}}$.
@@
This is [2, Theorem 5.1(2)]. See also [2, Theorem 1.2(2) or Proposition 4.8(2)] for essentially equivalent formulations of this theorem. A further discussion in category-theoretic context can be found in [15, Corollary 9.2 and Proposition 9.4]. ∎
@@
$\dotsb\mskip 1.5mu\relbar\joinrel\relbar\joinrel\rightarrow\mskip 1.5muC_{2}\mskip 1.5mu\relbar\joinrel\relbar\joinrel\rightarrow\mskip 1.5muC_{1}\mskip 1.5mu\relbar\joinrel\relbar\joinrel\rightarrow\mskip 1.5muC_{0}\overset{p}{\mskip 1.5mu\relbar\joinrel\relbar\joinrel\rightarrow\mskip 1.5mu}M\mskip 1.5mu\relbar\joinrel\relbar\joinrel\rightarrow\mskip 1.5mu0$
@@
This is [11, Corollary 1.6.4]. The point is that any $A$-module can be embedded into a cotorsion $A$-module (in fact, all injective $A$-modules are cotorsion by definition). Therefore, any $A$-module $M$ has a coresolution (right resolution) by cotorsion $A$-modules, i. e., there is an exact sequence of $A$-modules
$0\mskip 1.5mu\relbar\joinrel\relbar\joinrel\rightarrow\mskip 1.5muM\overset{i}{\mskip 1.5mu\relbar\joinrel\relbar\joinrel\rightarrow\mskip 1.5mu}C^{0}\mskip 1.5mu\relbar\joinrel\relbar\joinrel\rightarrow\mskip 1.5muC^{1}\mskip 1.5mu\relbar\joinrel\relbar\joinrel\rightarrow\mskip 1.5muC^{2}\mskip 1.5mu\relbar\joinrel\relbar\joinrel\rightarrow\mskip 1.5mu\dotsb$
@@
$\dotsb\mskip 1.5mu\relbar\joinrel\relbar\joinrel\rightarrow\mskip 1.5muC_{2}\mskip 1.5mu\relbar\joinrel\relbar\joinrel\rightarrow\mskip 1.5muC_{1}\mskip 1.5mu\relbar\joinrel\relbar\joinrel\rightarrow\mskip 1.5muC_{0}\overset{ip}{\mskip 1.5mu\relbar\joinrel\relbar\joinrel\rightarrow\mskip 1.5mu}C^{0}\mskip 1.5mu\relbar\joinrel\relbar\joinrel\rightarrow\mskip 1.5muC^{1}\mskip 1.5mu\relbar\joinrel\relbar\joinrel\rightarrow\mskip 1.5muC^{2}\mskip 1.5mu\relbar\joinrel\relbar\joinrel\rightarrow\mskip 1.5mu\dotsb$
@@
This is [11, Lemma 1.3.4(a)] or [14, Lemma 1.6(a)]. One can use the $\operatorname{Ext}^{1}$-adjunction of change of rings as in [5, Proposition VI.4.1.3] (see Lemma 3.4 below). Alternatively, the following elementary argument works. A left $R$-module $C$ is cotorsion if and only if the functor $\operatorname{Hom}_{R}({-},C)$ takes every short exact sequence of flat left $R$-modules to a short exact sequence of abelian groups. Given a short exact sequence of flat left $R$-modules $0\longrightarrow H\longrightarrow G\longrightarrow F\longrightarrow 0$, we have a short exact sequence of flat left $A$-modules $0\longrightarrow A\otimes_{R}H\longrightarrow A\otimes_{R}G\longrightarrow A\otimes_{R}F\longrightarrow 0$. By assumption, the $A$-module $D$ is cotorsion, so the functor $\operatorname{Hom}_{A}({-},D)$ takes the latter short exact sequence of $A$-modules to a short exact sequence of abelian groups. It remains to recall the adjunction isomorphism of abelian groups $\operatorname{Hom}_{A}(A\otimes_{R}M,\>D)\simeq\operatorname{Hom}_{R}(M,D)$, which holds for any left $R$-module $M$. The assertion of the lemma is also a special case of the next Lemma 2.4. ∎
@@
This is [11, Lemma 1.3.3(a)] or [14, Lemma 1.5(a)]. The point is that the tensor product $K\otimes_{S}G$ is a flat left $R$-module for any flat left $S$-module $G$. Based on this observation, one can use the $\operatorname{Ext}^{1}$-adjunction of the tensor product and Hom functors as in [5, formula (4) in Section XVI.4], [11, formula (1.1) in Section 1.2], or [14, Lemma 1.1]. Alternatively, an elementary argument can be found in [14, proof of Lemma 1.5]. ∎
@@
Let $R\longrightarrow A$ be a homomorphism of associative rings making $A$ a finitely generated projective right $R$-module. Consider the $R$-AA-bimodule
@@
By Lemma 2.5, we have a left $A$-module isomorphism $A\otimes_{R}C\simeq\operatorname{Hom}_{R}(K,C)$, where $K=\operatorname{Hom}_{R^{\mathrm{op}}}(A,R)$. The $R$-AA-bimodule $K$ is finitely generated and projective, hence flat, as a left $R$-module. So Lemma 2.4 is applicable for the ring $S=A$. ∎
@@
The lemma claims that the complex (3) is acyclic. In fact, the complex (3) is contractible as a complex of left $R$-modules (but not as a complex of left AA-modules). The natural $R$-linear contracting homotopy $h$ is given by the formula
@@
Proceeding in this way, we prove by induction that the left $A$-module $A^{\otimes_{R}\,n}\otimes_{R}D$ (with the $A$-module structure coming from the left $A$-module structure on the leftmost tensor factor $A$) is cotorsion for all $n\geq 1$. Now Lemma 2.7 provides a (left) resolution of $D$ by cotorsion $A$-modules. By Corollary 2.2, we can conclude that $D$ is a cotorsion $A$-module, as desired. ∎